OpenStudy (anonymous):

what is the nth term of the arithmetic series 3+7+11+15+19 Can you please explain to me step by step? Thanks a ton!!

OpenStudy (anonymous):

You can see each term increases by 4. The first one is given by taking the term number, multiplying by 4, and subtracting 1. 1x4-1=3 The same for the rest. 2x-4-1=7 So n times 4 minus 1 gives the nth term

OpenStudy (anonymous):

that's 2x4

OpenStudy (nincompoop):

TELLING YOU STEP-BY-STEP WILL NOT HELP SO MUCH. YOU HAVE TO READ THIS http://www.regentsprep.org/Regents/math/algtrig/ATP2/ArithSeq.htm

OpenStudy (anonymous):

is there another way to get the nth term, rather then just looking at it and figuring it out?

OpenStudy (nincompoop):

WHEN I GIVE YOU SOMETHING TO READ, READ IT.

OpenStudy (dmnts):

Answer is: 3+4n When n=0 you get 3, n=1 => 7, n=2 => 11 and so on...

OpenStudy (anonymous):

the answer is 4n-1 nincompoop : from the link you gave me, #6 was very similar to my question but I don't think I can look at a problem and figure it out,since the problems will most likely be harder